# SuperDiff: Diffusion Models for Conditional Generation of Hypothetical New Families of Superconductors

### 31 Jan 2024 v1

@@ Abstract
Effective computational search holds great potential for aiding the discovery of High-Temperature Superconductors (HTSs), especially given the lack of systematic methods for their discovery. Recent progress has been made in this area with machine learning, especially with deep generative models, which have been able to outperform traditional manual searches at predicting new superconductors within existing superconductor families but have yet to be able to generate completely new families of superconductors. We address this limitation by implementing conditioning—a method to control the generation process—for our generative model and develop SuperDiff, a Denoising Diffusion Probabilistic Model (DDPM) with Iterative Latent Variable Refinement (ILVR) conditioning for HTS discovery—the first deep generative model for superconductor discovery with conditioning. With SuperDiff, by being able to control the generation process, we were able to computationally generate completely new families of hypothetical superconductors for the very first time. Given that SuperDiff also has relatively fast training and inference times, it has the potential to be a very powerful tool for accelerating the discovery of new superconductors and enhancing our understanding of them.
@@ I Introduction
Understanding this, more recently, computational techniques have been applied to assist researchers in the search for new high $T_{c}$ superconductors. Specifically, a number of works have applied machine learning to this search for superconductors. Although serving as very valuable tools in many respects, most of these attempts [3, 4, 5], have been limited to classification and regression models, which only search through existing databases and are not able to generate any new compounds. Only recently, with deep generative models applied to superconductor discovery, have new hypothetical superconductors not found in most popular compound datasets been generated [6, 7]. In Ref. [6], a Generative Adversarial Network (GAN) [8] was applied for unconditional high $T_{c}$ superconductor generation, and in Ref. [7], a Crystal Diffusion Variational Autoencoder (CDVAE) [9] was also applied for unconditional superconductor generation so that crystal structure could be accounted for. However, Ref. [7] was applied on a different dataset and focused on a different task of generating stoichiometric Bardeen–Cooper–Schrieffer (BCS) conventional superconductors [10] and so did not generate any superconductors with $T_{c}\gtrsim$ 20 K.
@@
New attempts at high $T_{c}$ superconductor discovery with generative models are not without limitations, however. Most notably, although past models have been able to successfully generate new superconductors within existing superconductor families, they have not been able to generate completely new families of superconductors, which would be particularly desirable. This is because they are only unconditional models, which learn only the training dataset distribution. As unconditional models, the generation process of these models cannot be controlled. In other words, past models lack conditioning functionality—a method for controlling the generation process, that, in this context, means given an example superconductor, having the model generate similar superconductors, ideally by interpolating between the example and what the model has learned from the training dataset. With conditioning, the possibility of generating new families of superconductors can be opened, and researchers can be given control over the generation process. This can be especially useful for researchers looking to find only specific types of superconductors or expand on their own new discoveries. Moreover, Ref. [6] also struggled at generating unique (distinct from others in the given generated set) pnictides because of the small number of pnictides in SuperCon, the training dataset.
@@
To address these limitations, in this work, we implement a Denoising Diffusion Probabilistic Model (DDPM) [11, 12] for superconductor generation as our unconditional model and further implement conditioning with the Iterative Latent Variable Refinement (ILVR) [13] extension to DDPM, which allows for one-shot generation without additional training. With conditioning, we hope to be able to generate new families of superconductors, as identified by the clustering analysis proposed in Ref. [14], by experimenting with feeding the model different reference superconductors—this would mark a leap in the capabilities of computational searches for superconductors.
@@
We coin this first approach to conditionally generating new superconductors as “SuperDiff”. SuperDiff is also the first DDPM for superconductor generation. With SuperDiff, we aim to resolve the issues found in past works as a result of the small pnictide training dataset with the unconditional DDPM and, as our main focus, explore how the conditional DDPM can adapt to new information to generate new families of superconductors. This paper is structured as follows: in Section II, we give an overview of the underlying models behind SuperDiff and detail how they were applied; in Section III, we present the main results of the study, most notably a selection of new families of superconductors generated by SuperDiff—a first for the computational search for superconductors; and in Section IV, we summarize our contributions and most important findings.
@@ II.2 Denoising Diffusion Probabilistic Model
Four different versions of the DDPM were trained on SuperCon: one for cuprates, one for pnictides, one for others, and one for all classes (“everything”). The training datasets for each version of the DDPM were randomly split into training and validation sets in an approximately $95\%-5\%$ proportion. Training curves for all versions of the DDPM were able to converge and stabilize after around 50 epochs, and each version of the DDPM was trained for between 50 and 100 epochs, depending on the approximate lowest validation loss.
For all versions of the DDPM, NAdam [24] was chosen as the optimizer, and provided satisfactory results. Moreover, like in Ref. [11], $T$ was set to 1000 and the U-Net [25] neural network architecture was used for $\bm{\epsilon}_{\theta}$ (for this work, a 1D U-Net was used as opposed to the 2D U-Net used for images).
@@ IV Conclusion
With the lack of a systematic approach, the discovery of new high $T_{c}$ superconductors has long depended on material scientists’ serendipity. Recently, machine learning has been applied to this field to help assist scientists, but past works still lacked many key capabilities, for instance, the ability to computationally find new families of superconductors. Moreover, recent generative model approaches applied to this field also lacked methods of controlling the generation process [6, 7].

### 02 Feb 2024 v2

@@ Abstract
Effective computational search holds great potential for aiding the discovery of High-Temperature Superconductors (HTSs), especially given the lack of systematic methods for their discovery. Recent progress has been made in this area with machine learning, especially with deep generative models, which have been able to outperform traditional manual searches at predicting new superconductors within existing superconductor families but have yet to be able to generate completely new families of superconductors. We address this limitation by implementing conditioning—a method to control the generation process—for our generative model and develop SuperDiff, a Denoising Diffusion Probabilistic Model (DDPM) with Iterative Latent Variable Refinement (ILVR) conditioning for HTS discovery—the first deep generative model for superconductor discovery with conditioning on reference compounds. With SuperDiff, by being able to control the generation process, we were able to computationally generate completely new families of hypothetical superconductors for the very first time. Given that SuperDiff also has relatively fast training and inference times, it has the potential to be a very powerful tool for accelerating the discovery of new superconductors and enhancing our understanding of them.
@@ I Introduction
Understanding this, more recently, computational techniques have been applied to assist researchers in the search for new high $T_{c}$ superconductors. Specifically, a number of works have applied machine learning to this search for superconductors. Although serving as very valuable tools in many respects, most of these attempts [3, 4, 5], have been limited to classification and regression models, which only search through existing databases and are not able to generate any new compounds. Only recently, with deep generative models applied to superconductor discovery, have new hypothetical superconductors not found in most popular compound datasets been generated [6, 7, 8]. In Ref. [6], a Generative Adversarial Network (GAN) [9] was applied for unconditional high $T_{c}$ superconductor generation, and in Ref. [7], a Crystal Diffusion Variational Autoencoder (CDVAE) [10] was also applied for unconditional superconductor generation so that crystal structure could be accounted for. However, Ref. [7] was applied on a different dataset and focused on a different task of generating stoichiometric Bardeen–Cooper–Schrieffer (BCS) conventional superconductors [11] and so did not generate any superconductors with $T_{c}\gtrsim$ 20 K.
@@
New attempts at high $T_{c}$ superconductor discovery with generative models are not without limitations, however. Most notably, although past models have been able to successfully generate new superconductors within existing superconductor families, they have not been able to generate completely new families of superconductors, which would be particularly desirable. This is because they are only unconditional models, which learn only the training dataset distribution. As unconditional models, the generation process of these models cannot be controlled. In other words, past models lack conditioning functionality—a method for controlling the generation process, that, in this context, means giving an example superconductor, the reference compound, and having the model generate similar superconductors, ideally by interpolating between the example and what the model has learned from the training dataset. With conditioning, the possibility of generating new families of superconductors can be opened, and researchers can be given control over the generation process. This can be especially useful for researchers looking to find only specific types of superconductors or expand on their own new discoveries. Parallel to our work, Ref. [8] also applied a diffusion model for high $T_{c}$ superconductor discovery; however, their model was, like previous GANs, greatly limited by its lack of support for conditional generation with reference compounds—which is our main focus. Thus, their diffusion model shared with previous models the major limitation of being unable to generate any new families of superconductors—essentially, their work was only recreating the performance of the GAN in Ref. [6] but with a diffusion model instead and added $T_{c}$ label control only. Once again, we note that, in this work, we consider “conditioning” to mean conditioning the model on reference compounds only, as only this allows for the controlled generation of known and new families of superconductors. Moreover, Ref. [6] also struggled at generating unique (distinct from others in the given generated set) pnictides because of the small number of pnictides in SuperCon, the training dataset.
@@
To resolve these limitations, in this work, we implement a Denoising Diffusion Probabilistic Model (DDPM) [12, 13] for superconductor generation as our unconditional model and further implement conditioning with the Iterative Latent Variable Refinement (ILVR) [14] extension to DDPM, which allows for one-shot generation without additional training. With conditioning, we hope to be able to generate new families of superconductors for the first time, as identified by the clustering analysis proposed in Ref. [15], by experimenting with feeding the model different reference superconductors—this would mark a leap in the capabilities of computational searches for superconductors.
@@
We coin this first approach to conditionally generating new superconductors with reference compounds as “SuperDiff”. With SuperDiff, we aim to resolve the issues found in past works as a result of the small pnictide training dataset with the unconditional DDPM and, as our main focus, explore how the conditional DDPM can adapt to new information to generate completely new families of superconductors for the first time. This paper is structured as follows: in Section II, we give an overview of the underlying models behind SuperDiff and detail how they were applied; in Section III, we present the main results of the study, most notably a selection of new families of superconductors generated by SuperDiff—a first for the computational search for superconductors; and in Section IV, we summarize our contributions and most important findings.
@@ II.2 Denoising Diffusion Probabilistic Model
Four different versions of the DDPM were trained on SuperCon: one for cuprates, one for pnictides, one for others, and one for all classes (“everything”). The training datasets for each version of the DDPM were randomly split into training and validation sets in an approximately $95\%-5\%$ proportion. Training curves for all versions of the DDPM were able to converge and stabilize after around 50 epochs, and each version of the DDPM was trained for between 50 and 100 epochs, depending on the approximate lowest validation loss. For all versions of the DDPM, NAdam [25] was chosen as the optimizer, and provided satisfactory results. Moreover, like in Ref. [12], $T$ was set to 1000 and the U-Net [26] neural network architecture was used for $\bm{\epsilon}_{\theta}$ (for this work, a 1D U-Net was used as opposed to the 2D U-Net used for images).
@@ IV Conclusion
With the lack of a systematic approach, the discovery of new high $T_{c}$ superconductors has long depended on material scientists’ serendipity. Recently, machine learning has been applied to this field to help assist scientists, but past works still lacked many key capabilities, for instance, the ability to computationally find new families of superconductors. Moreover, recent generative model approaches applied to this field also lacked methods of controlling the generation process by incorporating information from reference compounds [6, 7, 8].